OpenStudy (anonymous):

if an object is dropped from a height of 85 feet the function h(t)=-16^2+85 gives the height of the object after t second . Approximately how long will it take for the object to hit the ground?

OpenStudy (ybarrap):

Solve for h(t)=85 for t. Does this make sense?

OpenStudy (anonymous):

i have no idea how to do these

OpenStudy (ybarrap):

Is the equation $$ h(t)=-16t^2+85 $$ ?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

Let's solve h(t) for t: $$ h(t)=-16t^2+85\\ -16t^2=h(t)-85\\ 16t^2=-h(t)+85\\ t^2=\cfrac{-h(t)+85}{16}\\ t=\sqrt{\cfrac{-h(t)+85}{16}} $$ Do you know what to do from here?

OpenStudy (ybarrap):

?

OpenStudy (ipwnbunnies):

I think it'd be easier to go back to ybarrap"s first suggestion. Set h(t) = 0, then solve for t. When h(t) = 0, that means the object is at ground level. The problem becomes a simple two-step algebra problem.

OpenStudy (ipwnbunnies):

\[h(t) = 0 = -16t^{2} + 85\]